show that a-b and a+b are consecutive terms of an ap

Grade:9

2 Answers

Neha
5 years ago
We need to show that a-b &a+b are consecutive terms of an ap
Now, let a1=a-b
& Common difference =b
Then a2 =a1+ common difference
a2=a-b+b
a2=a
a3 = a2 +common difference
a3= a+ b
Hence from above result 
we can prove that a-b& a+b are consecutive terms of an ap.
Hope it helps.
 
Aditya Gupta
5 years ago
The question is dumb and wrong, because any two terms can never decide the type of sequence. Any two random terms are not only in Ap, they are also in gp, hp, agp etc. So check your question and post again